You are walking around your neighborhood and you see a child on top of a roof of a building kick a soccer ball. The soccer ball is kicked at 33° from the edge of the building with an initial velocity of 21 m/s and lands 65 meters away from the wall. How tall, in meters, is the building that the child is standing on?
1
Expert's answer
2020-02-20T09:12:47-0500

Answer:-

Resolve the initial velocity X-component

"V_{x}=V_{0}cos\theta=21\times cos33\degree=17.62 \frac m s"

Resolve the initial velocity Y-component

"V_{y}=V_{0}sin\theta=21\times sin33\degree=11.44 \frac m s"

Time taken to travel 65 m along X-axis is.

"\to s=v_{x}t\\\rightarrow t=\frac{s}{v_{x}}=\frac{65}{17.62}\ =3.68 sec"


Total flight time of the ball is 3.68 sec

Now,

Find the height of building as follows.


"S_{y}-S_{_{yo}}=V_{y}t-\frac{1}{2}gt^{2}\\"

Here,

"\because" Sy=0 (final height)

Vy= 11.44 m/s (Initial velocity)

Find

Syo= ? (Initial height)


"0-S_{yo}=v_{y}t-\frac{1}{2}gt^{2}\\"

Now,

"S_{yo}=\left ( \frac{1}{2}\times 9.81\times (3.68)^{2} \right )-\left ( 11.44\times 3.68 \right )\\"


"S_{yo}=24.32m"


So,

Answer:-

The child is on a roof that is 24.32 m tall